OpenStudy (anonymous):

Given the following two points, (a) find the distance between them, (b) find the midpoint, and (c) submit the graph. Please show all of your work. (6, 1), (8, 0)

OpenStudy (phi):

distance between the 2 points can be found using the pythagorean thm d= sqrt( (8-6)^2+ (0-1)^2) = sqrt(2*2 +1*1)= sqrt(5) mid point is the average of the x values, y values x= (6+8/)/2 = 7 y = (1+0)/2 = 0.5

OpenStudy (anonymous):

phi- is this the answer then?

OpenStudy (anonymous):

how do i find the graph?

OpenStudy (phi):

plotting a graph means drawing an x-axis horizontal, a y-axis vertical and marking off numbers: